# Non-Hermitian Hamiltonian Approach for Two-Dimensional Spectroscopy

### v1

@@ 2.1 Response Function Formalism
$N_{p}=\left(1+\left|\beta_{p}\right|^{2}\right)^{-1/2},$
with $\delta t_{p}$ ($p=a,b,c$) being time intervals of the probe pulses.
Straightforward, the final density matrix reads $\rho(t)=|\psi\left(t_{3},t_{2},t_{1}\right)\rangle\langle\psi\left(t_{3},t_{2},t_{1}\right)|$. Under the situation that the spacing of the energy levels are much wider than their linewidths, its non-diagonal elements $\rho_{eb}$ and $\rho_{be}$ emit signals near $\omega_{e}$ in the frequency domain, which are proportional to the polarization
$\vec{P}_{eb}(t_{3},t_{2},t_{1})=\rho_{eb}(t_{3},t_{2},t_{1})\vec{\mu}_{be}+\rm c.c.$ (39)
Here, $\vec{\mu}_{be}$ is the transition electric dipole between $|e\rangle$ and $|b\rangle$. On account of the phase-matching condition, the polarization $\vec{P}_{\rm RP}(t_{3},t_{2},t_{1})$ in the RP direction along $\vec{k}_{s}=-\vec{k}_{a}+\vec{k}_{b}+\vec{k}_{c}$ is selected. Furthermore, by double Fourier transforms with respect to $t_{1}$ and $t_{3}$, the RP signal of the 2DS reads
@@
$\vec{P}_{\rm RP}(\omega_{3},t_{2},$ $\omega_{1})=e^{-i\omega_{e}t_{f}}N_{a}^{2}N_{b}N_{c}\beta_{a}^{*}\times$
@@
Here, $t_{f}=\sum_{p}\delta t_{p}$, and $\exp(-i\omega_{e}t_{c})$ are absorbed into $G_{ijkl}(t)$ and $\tilde{G}_{ijkl}(\omega)$. The quasi-Green functions and their Fourier transform for the NHH method are written as
@@ 3 Results and Discussion
In this section, we compare the 2DS by the two methods with the same parameters, i.e., the ground-state energy $\omega_{b}=0$, the excited-state energy $\omega_{e}/2\pi=4.33~{}\rm THz$, and the metastable-state energy $\omega_{c}/2\pi=1~{}\rm MHz$ 45, 46, 47. The Rabi frequencies and durations of the probe pulses are $\Omega_{be}/2\pi=50~{}\rm MHz$ and $\delta t_{p}=0.5~{}\rm ns$, with a bandwidth of $2\pi\times 1.8~{}\rm GHz$, while the Rabi frequency of the control field is $\Omega_{ec}=2~{}\rm MHz$ 48. The dissipation rates are $\Gamma_{1}/2\pi=1~{}\rm kHz$ and $\Gamma_{2}/2\pi=0.03~{}\rm kHz$, and the pure-dephasing rates of all states are homogeneously $\gamma_{b}^{\left(0\right)}/2\pi=\gamma_{e}^{\left(0\right)}/2\pi=\gamma_{c}^{\left(0\right)}/2\pi=0.1~{}\rm MHz$ 49, 50.
@@
Figure 4: Comparison of the Green functions by the RF ($\mathcal{\tilde{G}}_{ijkl}$) and the NHH ($\tilde{G}_{ijkl}$) approaches. (a) $\mathcal{\tilde{G}}_{bebe}$ vs $i\tilde{G}_{bebe}$, and $i\tilde{G}_{bcbe}$ are respectively denoted by the orange solid, dark-green circles, light-green dash lines, (b) $\mathcal{\tilde{G}}_{ebeb}$ vs $-i\tilde{G}_{ebeb}$, and $-i\tilde{G}_{ebcb}$ are respectively denoted by the purple solid, dark-blue circles, light-blue dash lines.
@@
The comparison between the Green function $\mathcal{\tilde{G}}_{ijkl}(\omega)$ and the quasi-Green functions $\tilde{G}_{ijkl}(\omega)$ are shown in Fig. 4. The additional factors ±$\pm i$ before $\tilde{G}_{ijkl}(\omega)$ are included due to the different definitions of the Fourier transforms in the RF and the NHH approaches 29. The curves of Im($\mathcal{\tilde{G}}_{bebe}$) and Im($i\tilde{G}_{bebe}$) are identical, and so are Im($\mathcal{\tilde{G}}_{ebeb}$) and Im($-i\tilde{G}_{ebeb}$). Concurrently, the curve of Im($i\tilde{G}_{bcbe}$) coincides with Im($\mathcal{\tilde{G}}_{bebe}$) and Im($i\tilde{G}_{bebe}$) when $\omega-\omega_{e}<0$, and differs from them by a sign when $\omega-\omega_{e}>0$. The same behavior can be observed in the curves of Im($-i\tilde{G}_{ebcb}$) vs Im($\mathcal{\tilde{G}}_{ebeb}$) and Im($-i\tilde{G}_{ebeb}$). The imaginary parts of the Green functions for $\omega_{1}$ dominate the absorption of probe pulses, with higher positive values representing greater absorption and lower negative values representing more emission. The opposite sign of the imaginary parts of the Green functions for $\omega_{3}$ have the same influence. The control field causes the original absorption peak at $\omega_{e}$ to be split into two peaks with approximately-symmetrical locations around $\omega_{e}$, just as EIT happens. Furthermore, the emergent evolution paths of $\tilde{G}_{bcbe}$ and $\tilde{G}_{ebcb}$ introduce emission peaks. Consequently, the splitting determines the location of peaks on 2DS, and the influence of the emission peaks will be discussed later.
@@
Figure 5: Population dynamics by (a) the RF formalism, (b) the NHH method and (c) the Green function $G_{ijkl}(t)$ for the coherences vs time. The dark-green solid and dark-blue dash-dotted lines denote $\mathcal{G}_{bbbb}$ ($G_{bbbb}$) and $\mathcal{G}_{eeee}$ ($G_{eeee}$), respectively. The light-blue dotted and dark-red dashed lines respectively denote $\mathcal{G}_{bbee}$ and $\mathcal{G}_{eebb}$, while the light-green solid, orange dashed, and purple dotted lines respectively denote $\textrm{Im}(G_{ceee})$, $\textrm{Im}(G_{eeec})$ and $G_{ceec}$. Here, $G_{ceee}$ and $G_{eeec}$ are purely imaginary. The inset in (a) shows the long-time regime, and the black circles in the first two subfigures label $t_{2}=0.24~{}\mu$s and $t_{2}=0.50~{}\mu$s, respectively.
@@
There is a phase shift of $\pi$ between the oscillations of $G_{eeee}$ and $G_{ceec}$. When $G_{eeee}=0$, $G_{ceec}$ reaches its maximum values, indicating that the maximum population transfer corresponds to the maximum coherence. In addition, the imaginary parts of $G_{ceee}$ and $G_{eeec}$ exhibit inverse-synchronized damping oscillations. They oscillate around zero, with a phase shift of $\pi$ between them as well. However, they have no direct effect on the final 2DS. Moreover, although the oscillation behaviors by the two methods exhibit similarity, they tend to reach different steady states in long-time limit, e.g., $\Gamma_{1}/(\Gamma_{1}+\Gamma_{2})$ for $\mathcal{G}_{bbbb}$ vs zero for $G_{bbbb}$, according to Eqs. (19) and (28). This remarkable discrepancy is due to the fact that the Hamiltonian in the NHH method is not Hermitian, i.e., $H^{\dagger}\neq H$, so the conservation of the trace is violated and thus the detailed balance breaks down.
@@
Figure 7: Comparison of the 2DS for the two methods. The first (second) row correspond to the RF formalism (the NHH method). The third row is by the RF formalism with Liouville paths of the same as the NHH method in Fig. 3. The 2DS are plotted when (a,e,i) $t_{2}=0$ $\rm\mu s$, (b,f,j) $t_{2}=0.24$ $\rm\mu s$, (c,g,k) $t_{2}=0.50$ $\rm\mu s$, (d,h,l) $t_{2}=10^{3}$ $\rm\mu s$.
@@
For further comparison of the two approaches, representative points are selected according to their population dynamics, i.e., $t_{2}=0,~{}0.24,~{}0.5,~{}10^{3}$ $\rm\mu s$, and the 2DS for each method at these population times are plotted in Fig. 7 (a) through (h). The first three $t_{2}$ values correspond to the initial state, the first valley and the first peak of the population dynamics, respectively, for both the RF formalism and the NHH method. The last $t_{2}$ is selected for observing results at the steady state. In both methods, the introduction of the control field results in splitting of the homogeneously-broadened peak into four smaller peaks on 2DS 44. The RF formalism and the NHH method, exhibiting identical peak positions, indicate that the spacings between peaks are identical. These spacings between the adjacent diagonal and non-diagonal peaks equal to $2.005\times 2\pi~{}\rm MHz$, which are approximately the splitting width $2\times 2\pi~{}\rm MHz$ of the Green functions according to Fig. 4.
@@
The RP signals of the two methods are respectively normalized by their maximum values at $t_{2}=0~{}\rm\mu s$, thereby obtaining relative signals for the sake of simplicity. At $t_{2}=0~{}\rm\mu s$, 2DS of the two methods are close to each other. Both methods yield the highest four peaks, the relative heights of one peak and its counterpart from the other method are approximately identical, with a variance at the scale of $10^{-4}$. At $t_{2}=10^{3}~{}\rm\mu s$, the peaks of the RF formalism retain their scale because the trace is conserved. The peaks of the NHH method, however, undergo an overall attenuation as $t_{2}$ increases, since its trace is not conserved. The peak heights rapidly attenuate in the first few seconds, and are of the order of $10^{-273}$ different from the initial state at $t_{2}=10^{3}~{}\rm\mu s$.
@@
Figure 8: Comparison of the NHH method’s partial 2DS with different evolving paths in $t_{1}$ and $t_{3}$. The first (second) row correspond to the summation of the paths evolving from $|b\rangle\langle e|$ to $|b\rangle\langle e|$ ($|b\rangle\langle e|$ to $|b\rangle\langle c|$) in $t_{1}$, while evolving from $|e\rangle\langle b|$ to $|e\rangle\langle b|$ ($|c\rangle\langle b|$ to $|e\rangle\langle b|$) in $t_{3}$. The 2DS are plotted when (a,c) $t_{2}=0.24$ $\rm\mu s$, (b,d) $t_{2}=0.50$ $\rm\mu s$.
@@
The four peaks of the RF formalism synchronously become lower at $t_{2}=~{}0.24~{}\rm\mu s$, then raise at $t_{2}=~{}0.5~{}\rm\mu s$. They exhibit synchronous oscillating behavior, which is dominated by $\mathcal{G}_{eeee}$. The NHH method, on the other hand, has little variance on the diagonal peaks despite attenuation, while the non-diagonal peaks continue to oscillate. According to the Liouville paths of each term in $t_{1}$ and $t_{3}$, the RP signal of the NHH method is divided into three parts. The paths of Fig. 3(a) and (b) are added together as the part that experiences $|b\rangle\langle e|$ evolving to $|b\rangle\langle e|$ during $t_{1}$, while $|e\rangle\langle b|$ evolving to $|e\rangle\langle b|$ during $t_{3}$. Figure 3(c) and (d) are added together and have little effect on the final 2DS. Figure 3(e) is classified into a separate part for it experiences time evolution from $|b\rangle\langle e|$ to $|b\rangle\langle c|$ in $t_{1}$ and $|c\rangle\langle b|$ to $|e\rangle\langle b|$ in $t_{3}$. The 2DS of the first and the third parts are demonstrated in Fig. 8. In Fig. 8(c), the non-diagonal peaks are negative, which are caused by the emission peaks of the quasi-Green function in Fig. 4. The remaining positive peaks at $t_{2}=0.24$ $\rm\mu s$ are shallow, but are summed up to be the same scale in accordance with the peak heights at $t_{2}=0$ $\rm\mu s$. At $t_{2}=0.50$ $\rm\mu s$, on the other hand, Fig. 8(d) gives little contribution and the final 2DS is dominated by Fig. 8(b). As a result, Fig. 8(b) is almost the same as Fig. 7(g). These 2DS are consistent with the population dynamics. In other words, having less population on the excited state $|e\rangle$ despite decay is the feature of stronger coherence between the energy levels $e$ and $c$, as well as more contribution from the Liouville path Fig. 3(e).
@@
Finally, for a better comparison of the two approaches, we recalculate the 2DS by the RF formalism with Liouville paths the same as the NHH method in Fig. 3. The RP signal is written as
@@
$+\frac{1}{2}e^{-\gamma_{ec}^{+}t},$ (53)
and the 2DS are demonstrated in Fig. 7 (i) through (l). It is shown that when the RF formalism and the NHH method are applied using the same Liouville paths, the resulting 2DS exhibit quite similar behaviors. The only discrepancy of the 2DS occurs significantly in the long-time limit, where the peak heights by the NHH method markedly decrease as compared to the counterparts by the RF formalism. These results reveal that the two methods are mostly consistent when having the identical Liouville paths.

### v2

@@ 2.1 Response Function Formalism
$N_{p}=\left(1+\left|\beta_{p}\right|^{2}\right)^{-1/2}.$
Straightforward, the final density matrix reads $\rho^{I}(t)=|\psi^{I}\left(t_{3},t_{2},t_{1}\right)\rangle\langle\psi^{I}\left(t_{3},t_{2},t_{1}\right)|$, and the total polarization is $P(t)=\langle\mu^{I}(t)\rho^{I}(t)\rangle$. Under the situation that the spacing of the energy levels are much wider than their linewidths, its non-diagonal elements $\rho^{I}_{eb}$ and $\rho^{I}_{be}$ emit signals near $\omega_{e}$ in the frequency domain, which are proportional to the polarization
$\vec{P}_{eb}(t_{3},t_{2},t_{1})=\rho_{eb}^{I}(t_{3},t_{2},t_{1}){\mu}^{I}_{be}(t_{1}+t_{2}+t_{3}+\delta t_{a}+\delta t_{b}+\delta t_{c})+\rm c.c.$ (39)
Here, ${\mu}^{I}_{be}(t)$ is the component corresponding to the transition from $|e\rangle$ to $|b\rangle$ of the electric-dipole operator ${\mu}^{I}(t)$. On account of the phase-matching condition, the polarization $\vec{P}_{\rm RP}(t_{3},t_{2},t_{1})$ in the RP direction along $\vec{k}_{s}=-\vec{k}_{a}+\vec{k}_{b}+\vec{k}_{c}$ is selected. Furthermore, by double Fourier transforms with respect to $t_{1}$ and $t_{3}$, the RP signal of the 2DCS reads
@@
$\vec{P}_{\rm RP}(\omega_{3},t_{2},$ $\omega_{1})=N_{a}^{2}N_{b}N_{c}\beta_{a}^{*}\times$
@@
Here, $\exp(-i\omega_{e}t_{c})$ are absorbed into $G_{ijkl}(t)$ and $\tilde{G}_{ijkl}(\omega)$. The quasi-Green functions and their Fourier transform $G_{ijkl}(t)$ and $\tilde{G}_{ijkl}(\omega)$ are proposed for the NHH method, and are written as
@@
Finally, a summary of the approximations used in the RF formalism and the NHH method is presented in Table 1, highlighting both the similarities and differences in their treatment. First of all, since the quantum-jump terms are neglected, the detailed balance can not be reproduced in the NHH method. Additionally, to simplify the calculations, the square-pulse approximation is adopted. However, more realistic pulse shapes can be considered if $H^{p}_{\textrm{int}}(t)$ is appropriately revised. More importantly, in the RF formalism, we treat the off-diagonal terms of the density matrix perturbatively to obtain the analytic solution. The more-accurate results may be achieved if this approximation is discarded.
Appr. Method RF formalism NHH method
Different perturbative $H_{\textrm{int}}^{p}$ square pulse appr.
semi-impulsive limit non-Hermitian $H$
perturbative $\rho_{ij}$
Same $H_{\textrm{p}}\simeq H_{0}+H_{\textrm{int}}^{p}$
$\nu_{c}=\omega_{ec}$
Table 1: Summary of approximations used in the two methods. “Appr.” is short for “Approximation”.
@@ 3 Results and Discussion
In this section, we compare the 2DCS of the two methods by the example of 1,2-propanediol gas at room temperature. Three states under investigation are selected with $|b\rangle=|v_{g}\rangle|0_{0,0,0}\rangle$, $|e\rangle=|v_{e}\rangle|1_{1,1,1}\rangle$ and $|c\rangle=|v_{e}\rangle|1_{0,1,0}\rangle$. $|v_{g}\rangle$ ($|v_{e}\rangle$) marks the vibrational ground (first-excited) state, and $|J_{K_{a},K_{c},M}\rangle$ marks the rotational states. The transition frequencies of each states are respectively $\omega_{b}=0$, $\omega_{e}/2\pi=4.33~{}\rm THz$, and $\omega_{c}/2\pi=4.32~{}\rm THz$ 71, 72, 73, 64. The Rabi frequencies and durations of the probe pulses are $\Omega_{be}/2\pi=50~{}\rm MHz$ and $\delta t_{p}=0.5~{}\rm ns$, with a bandwidth of $2\pi\times 1.8~{}\rm GHz<\omega_{ec}$, while the Rabi frequency of the control field is $\Omega_{ec}/2\pi=2~{}\rm MHz$ 64. The dissipation rates are $\Gamma_{1}/2\pi=1~{}\rm kHz$ and $\Gamma_{2}/2\pi=0.03~{}\rm kHz$, and the pure-dephasing rates of all states are homogeneously $\gamma_{b}^{\left(0\right)}/2\pi=\gamma_{e}^{\left(0\right)}/2\pi=\gamma_{c}^{\left(0\right)}/2\pi=0.1~{}\rm MHz$ 74, 75.
@@
Figure 4: Comparison of the Green functions by the RF ($\mathcal{\tilde{G}}_{ijkl}$) and the NHH ($\tilde{G}_{ijkl}$) approaches. (a) $\mathcal{\tilde{G}}_{bebe}$ vs $i\tilde{G}_{bebe}$, and $i\tilde{G}_{bcbe}$ are respectively denoted by the orange solid, dark-green circles, light-green dashed lines, (b) $\mathcal{\tilde{G}}_{ebeb}$ vs $-i\tilde{G}_{ebeb}$, and $-i\tilde{G}_{ebcb}$ are respectively denoted by the purple solid, dark-blue circles, light-blue dashed lines.
@@
The comparison between the Green function $\mathcal{\tilde{G}}_{ijkl}(\omega)$ and the quasi-Green functions $\tilde{G}_{ijkl}(\omega)$ are shown in Fig. 4. The additional factors ±$\pm i$ before $\tilde{G}_{ijkl}(\omega)$ are included due to the different definitions of the Fourier transforms in the RF and the NHH approaches 41. The curves of Im($\mathcal{\tilde{G}}_{bebe}$) and Im($i\tilde{G}_{bebe}$) are identical, and so are Im($\mathcal{\tilde{G}}_{ebeb}$) and Im($-i\tilde{G}_{ebeb}$). Concurrently, the curve of Im($i\tilde{G}_{bcbe}$) coincides with Im($\mathcal{\tilde{G}}_{bebe}$) and Im($i\tilde{G}_{bebe}$) when $\omega-\omega_{e}<0$, and differs from them by a sign when $\omega-\omega_{e}>0$. The same behavior can be observed in the curves of Im($-i\tilde{G}_{ebcb}$) vs Im($\mathcal{\tilde{G}}_{ebeb}$) and Im($-i\tilde{G}_{ebeb}$). The imaginary parts of the Green functions for $\omega_{1}$ dominate the absorption of probe pulses, with higher positive values representing greater absorption and lower negative values representing more emission. The opposite sign of the imaginary parts of the Green functions for $\omega_{3}$ have the same influence. The control field causes the original absorption peak at $\omega_{e}$ to be split into two peaks with approximately-symmetrical locations around $\omega_{e}$, just as EIT happens. Furthermore, the emergent evolution paths of $\tilde{G}_{bcbe}$ and $\tilde{G}_{ebcb}$ introduce emission peaks. Consequently, the splitting determines the location of peaks on 2DCS, and the influence of the emission peaks will be discussed later.
@@
Figure 5: Population dynamics by (a) the RF formalism, (b) the NHH method and (c) the quasi-Green function $G_{ijkl}(t)$ for the coherences vs time. The dark-green solid and dark-blue dash-dotted lines denote $\mathcal{G}_{bbbb}$ ($G_{bbbb}$) and $\mathcal{G}_{eeee}$ ($G_{eeee}$), respectively. The light-blue dotted and dark-red dashed lines respectively denote $\mathcal{G}_{bbee}$ and $\mathcal{G}_{eebb}$, while the light-green solid, orange dashed, and purple dotted lines respectively denote $\textrm{Im}(G_{ceee})$, $\textrm{Im}(G_{eeec})$ and $G_{ceec}$. Here, $G_{ceee}$ and $G_{eeec}$ are purely imaginary. The inset in (a) shows the long-time regime, and the black circles in the first two subfigures label $t_{2}=0.24~{}\mu$s and $t_{2}=0.50~{}\mu$s, respectively.
@@
There is a phase shift of $\pi$ between the oscillations of $G_{eeee}$ and $G_{ceec}$. When $G_{eeee}=0$, $G_{ceec}$ reaches its maximum values, indicating that the maximum population transfer corresponds to the maximum coherence. In addition, the imaginary parts of $G_{ceee}$ and $G_{eeec}$ exhibit anti-synchronized damping oscillations. They oscillate around zero, with a phase shift of $\pi$ between them as well. However, they have no direct effect on the final 2DCS. Moreover, although the oscillation behaviors by the two methods exhibit similarity, they tend to reach different steady states in long-time limit, e.g., $\Gamma_{1}/(\Gamma_{1}+\Gamma_{2})$ for $\mathcal{G}_{bbbb}$ vs zero for $G_{bbbb}$, according to Eqs. (19) and (28). This remarkable discrepancy is due to the fact that the Hamiltonian in the NHH method is not Hermitian, i.e., $H^{\dagger}\neq H$, so the conservation of the trace is violated and thus the detailed balance breaks down.
@@
Figure 7: Comparison of the 2DCS for the two methods. The first (second) row correspond to the RF formalism (the NHH method). The third row is by the RF formalism with Liouville paths of the same as the NHH method in Fig. 3. The 2DCS are plotted when (a,e,i) $t_{2}=0$ $\rm\mu s$, (b,f,j) $t_{2}=0.24$ $\rm\mu s$, (c,g,k) $t_{2}=0.50$ $\rm\mu s$, (d,h,l) $t_{2}=10^{3}$ $\rm\mu s$.
@@
For further comparison of the two approaches, representative points are selected according to their population dynamics, i.e., $t_{2}=0,~{}0.24,~{}0.5,~{}10^{3}$ $\rm\mu s$, and the 2DCS for each method at these population times are plotted in Fig. 7 (a) through (h). The first three $t_{2}$ values correspond to the initial state, the first valley and the first peak of the population dynamics, respectively, for both the RF formalism and the NHH method. The last $t_{2}$ is selected for observing results at the steady state. In both methods, the introduction of the control field results in splitting of the homogeneously-broadened peak into four smaller peaks on 2DCS 63. The RF formalism and the NHH method, exhibiting identical peak positions, indicate that the spacings between peaks are identical. These spacings between the adjacent diagonal and non-diagonal peaks equal to $2.005\times 2\pi~{}\rm MHz$, which are approximately the splitting width $2\times 2\pi~{}\rm MHz$ of the Green functions according to Fig. 4.
Figure 8: The heights of diagonal and non-diagonal peaks are shown by (a) the RF formalism, (b) the NHH method, and (c) the RF formalism with the Liouville paths from the NHH method, as a function of $t_{2}$. The light-blue dashed and purple dotted lines represent the diagonal and non-diagonal peak heights, respectively.
@@
The RP signals of the two methods are respectively normalized by their maximum values at $t_{2}=0~{}\rm\mu s$, thereby obtaining relative signals for the sake of simplicity. At $t_{2}=0~{}\rm\mu s$, both methods yield the highest four peaks, and the relative heights for one peak and its counterpart from the other method are approximately identical, with a variance at the scale of $10^{-4}$. At $t_{2}=10^{3}~{}\rm\mu s$, the peaks of the RF formalism retain their scale because the trace is conserved. The peaks of the NHH method, however, undergo overall attenuation as $t_{2}$ increases, since its trace is not conserved. The peak heights rapidly attenuate in the first few microseconds. Even so, Fig. 7 (h), with an extremely small colorbar (which is not shown here), is similar to Fig. 7 (d). This indicates that, although the NHH method continues to lose its scale, it still preserves the relative information between peaks. Here, we have the second condition that should be satisfied when using the NHH method. This method is suitable for short-time dynamics and fails to observe the steady state in the long-time limit.
@@
Figure 9: Comparison of the NHH method’s partial 2DCS with different evolving paths in $t_{1}$ and $t_{3}$. The first (second) row correspond to the summation of the paths evolving from $|b\rangle\langle e|$ to $|b\rangle\langle e|$ ($|b\rangle\langle e|$ to $|b\rangle\langle c|$) in $t_{1}$, while evolving from $|e\rangle\langle b|$ to $|e\rangle\langle b|$ ($|c\rangle\langle b|$ to $|e\rangle\langle b|$) in $t_{3}$. The 2DCS are plotted when (a,c) $t_{2}=0.24$ $\rm\mu s$, (b,d) $t_{2}=0.50$ $\rm\mu s$.
The four peaks of the RF formalism synchronously become lower at $t_{2}=~{}0.24~{}\rm\mu s$, then raise at $t_{2}=~{}0.5~{}\rm\mu s$. They exhibit synchronous oscillating behavior, which is dominated by $\mathcal{G}_{eeee}$. The NHH method, on the other hand, has little variance on the diagonal peaks despite attenuation, while the off-diagonal peaks continue to oscillate. For better clarity, the heights of the diagonal peak at (-1.005,-1.005) MHz and the non-diagonal peak at (-1.005,0.995) MHz are plotted as functions of $t_{2}$, as shown in Fig. 8 (a) and (b).
@@
According to the Liouville paths of each term in $t_{1}$ and $t_{3}$, the RP signal of the NHH method is divided into three parts. The paths of Fig. 3(a) and (b) are added together as the part that experiences $|b\rangle\langle e|$ evolving to $|b\rangle\langle e|$ during $t_{1}$, while $|e\rangle\langle b|$ evolving to $|e\rangle\langle b|$ during $t_{3}$. Figure 3(c) and (d) are added together and have little effect on the final 2DCS. Figure 3(e) is classified into a separate part for it experiences time evolution from $|b\rangle\langle e|$ to $|b\rangle\langle c|$ in $t_{1}$ and $|c\rangle\langle b|$ to $|e\rangle\langle b|$ in $t_{3}$. The 2DCS of the first and the third parts are demonstrated in Fig. 9. In Fig. 9(c), the non-diagonal peaks are negative, which are caused by the emission peaks of the quasi-Green function in Fig. 4. The remaining positive peaks at $t_{2}=0.24$ $\rm\mu s$ are shallow, but are summed up to be the same scale in accordance with the peak heights at $t_{2}=0$ $\rm\mu s$. At $t_{2}=0.50$ $\rm\mu s$, on the other hand, Fig. 9(d) gives little contribution and the final 2DCS is dominated by Fig. 9(b). As a result, Fig. 9(b) is almost the same as Fig. 7(g). These 2DCS are consistent with the population dynamics. In other words, having less population on the excited state $|e\rangle$ despite decay is the feature of stronger coherence between the energy levels $e$ and $c$, as well as more contribution from the Liouville path Fig. 3(e).
@@
Next, for a better comparison of the two approaches, we recalculate the 2DCS by the RF formalism with Liouville paths the same as the NHH method in Fig. 3. The RP signal is written as
@@
$+\frac{1}{2}e^{-\gamma_{ec}^{+}t}.$ (53)
In this case, the 2DCS are shown in Fig. 7 (i) to (l). Meanwhile, the variance of heights of the diagonal and non-diagonal peaks is plotted in Fig. 8 (c). It is shown that when the RF formalism and the NHH method are applied using the same Liouville paths, the resulting 2DCS exhibit quite similar behaviors. The only discrepancy of the 2DCS occurs significantly in the long-time limit, where the peak heights by the NHH method markedly decrease as compared to the counterparts by the RF formalism. These results reveal that the two methods are mostly consistent when having the identical Liouville paths.
@@
Figure 10: (a) Schematic diagram of the six levels in RDC, which includes the ground vibrational state $|00\rangle$, one-quantum states $|a\rangle$ and $|s\rangle$, as well as two-quantum states $|2a\rangle$, $|2s\rangle$ and $|as\rangle$. The transition frequencies are respectively $\omega_{a,0}=2015~{}\textrm{cm}^{-1}$, $\omega_{s,0}=2084~{}\textrm{cm}^{-1}$, $\omega_{2a,a}=2001~{}\textrm{cm}^{-1}$, $\omega_{2s,s}=2073~{}\textrm{cm}^{-1}$, $\omega_{as,a}=2058~{}\textrm{cm}^{-1}$, $\omega_{as,s}=1989~{}\textrm{cm}^{-1}$, $\omega_{2a,s}=1932~{}\textrm{cm}^{-1}$, and $\omega_{2s,a}=2142~{}\textrm{cm}^{-1}$. 76, 77. (b) 2DCS of RDC dissolved in hexane by the NHH method. The horizontal and vertical lines mark the frequencies $\omega_{a,0}$ and $\omega_{s,0}$. The gap values are approximately $\Delta_{as}=26~{}\textrm{cm}^{-1}$, $\Delta_{a}=14~{}\textrm{cm}^{-1}$, $\Delta_{s}=11~{}\textrm{cm}^{-1}$ 76, 77.
Finally, we simulate the 2DCS of RDC dissolved in hexane using the NHH method. The system has already been experimentally measured by Khalil et al. 76. This system has six states, which include a ground vibrational state $|00\rangle$, two one-quantum states, i.e., $|a\rangle$, $|s\rangle$, and three two-quantum states, i.e., $|2a\rangle$, $|2s\rangle$, $|as\rangle$. The spectral bandwidth of the three incident pulses is wide enough to excite all the vibrational transitions. The energy levels and transitions are shown in Fig. 10(a), with dashed arrows representing forbidden transitions and solid arrows representing allowed transitions. The corresponding frequencies are as follows 76, 77, $\omega_{a,0}=2015~{}\textrm{cm}^{-1}$, $\omega_{s,0}=2084~{}\textrm{cm}^{-1}$, $\omega_{2a,a}=2001~{}\textrm{cm}^{-1}$, $\omega_{2s,s}=2073~{}\textrm{cm}^{-1}$, $\omega_{as,a}=2058~{}\textrm{cm}^{-1}$, $\omega_{as,s}=1989~{}\textrm{cm}^{-1}$, $\omega_{2a,s}=1932~{}\textrm{cm}^{-1}$, and $\omega_{2s,a}=2142~{}\textrm{cm}^{-1}$, with the total decoherence rate $\gamma=0.3~{}\textrm{cm}^{-1}$. The elements of the electric dipole operator, describing the intensity of each transition, are given by $\mu_{a,0}=1.05\mu_{s,0},\mu_{as,s}=\mu_{a,0},\mu_{as,a}=\mu_{s,0},\mu_{aa,a}=1.48\mu_{s,0},\mu_{ss,s}=1.41\mu_{s,0},\mu_{ss,a}=\mu_{aa,s}=0.13\mu_{s,0}$ 78. Meanwhile, $t_{1}$ and $t_{3}$ vary from 0 to 5 ps, with a time step of $5~{}\rm{fs}$, and $t_{2}$ is set to 0. The pulse durations are neglected, as they contribute to the final results by an overall parameter.
The purely-absorptive spectrum is obtained by calculating the rephasing and non-rephasing signals. Zero padding is applied to expand the data length to 5000 points per dimension before performing the Fourier transform. After performing the double Fourier transforms, the real parts of the rephasing and non-rephasing signals are taken. The rephasing signal is then reflected against the $\it{y}$-axis and summed with the non-rephasing signal. As a result, the 2DCS for absorptive signals is shown in Fig. 10(b).
Our simulation qualitatively reproduces the experimental features demonstrated by Khalil et al., including the number, positions, signs of the peaks in 2DCS 76. The system can only be excited to the one-quantum states by the first pulse. As a result, the peaks appear along the $\omega_{1}$ axis only when $\omega_{1}=\omega_{a,0},\omega_{s,0}$. Along the $\omega_{3}$ axis, on the other hand, there are 6 possible frequencies at which peaks appear. When $\omega_{3}=\omega_{a,0},\omega_{s,0}$, formed by the excitation of the third pulse, the peaks have positive signs. When $\omega_{3}=\omega_{2a,a},\omega_{2s,s},\omega_{as,a},\omega_{as,s}$, the peaks have negative signs indicating the radiations from the two-quantum states. The gaps between the positive peaks and the negative peaks are thus equal to the frequency differences between the radiations and fundamental transitions. These gaps, marked by double-headed arrows on the plot, are approximately $\Delta_{as}=26~{}\textrm{cm}^{-1}$, $\Delta_{a}=14~{}\textrm{cm}^{-1}$, $\Delta_{s}=11~{}\textrm{cm}^{-1}$, which are equal to the frequency differences 76. Considering the well-fitted result, we believe the NHH method is a reliable theoretical approach for simulating the 2DCS.